OpenStudy (anonymous):

In a high school, the ratio of male to female students is 8:5. When 400 senior males to females are removed and 300 senior females are removed, the ration is 8:3. The number of students in the whole school is?

OpenStudy (anonymous):

Males = 8X Females = 5X \[\frac{ 8X-400 }{ 5X-300 }= \frac{8}{3}\] X = 75 Total = Males + Females = 8X + 5X = 13X = 13*75 = 975

OpenStudy (mathstudent55):

Let's say there are x males and y females. Then x/y = 8/5 When you remove the students, the ratio becomes: (x - 400)/(y - 300) = 8/3 Cross multiply both equations: 5x = 8y 3(x - 400) = 8(y - 300) 5x = 8y 3x - 1200 = 8y - 2400 Now subtract bottom eq from top eq: 2x + 1200 = 2400 2x = 1200 x = 600 5x = 8y 5(600) = 8y 8y = 3000 y = 375 Answer: 600 males + 375 females = 975 students

OpenStudy (anonymous):

let there be 8x and 5x males and females respectively then by question we have (8x-400)/(5x-300)=8/3 or 8(x-50)/(5/(x-60)=8/3 or 3(x-50)=5(x-60) 3x-150=5x-300 2x =150 x=75 total number of students is 8x+5x=13x =13*75 =975